In a type of cancer treatment called the thermotherapy, temperature as high as 113F are used to destroy cancer cells. What is that temperature in C and kelvin?
Answer #1

Celsius :(Fahrenheit _ 32)×5÷9

=(113_32)×5÷9

=81×5÷9

=405÷9

=45

The temperature in Celsius is 45°C

Kelvin:(F_32)×5÷9+273.15

=(113_32)×5÷9+273.15

=81×5÷9+273.15

=45+273.15

=318.15K

The temperature in Kelvin is 318.15K